About 3% of children in a country are allergic to peanuts. Choose three children at random and let the random variable X be the number in this sample who are allergic to peanuts. The possible values X can take are 0, 1, 2, and 3. What is the conditional probability that exactly two of the children will be allergic to peanuts, given that at least one of the three children suffers from this allergy? (Round your answer to four decimal places.)
Answer by Boreal(15235)

probability from 3
0: .97^3=0.91267
1: 3(.97^2)(0.03)=0.08468
2: 3(.97)(.03^2)=0.00262
3: .03^3=0.00003
given that at least 1 has it (prob. 0.08733) probability exactly 2 have it is 0.00262/0.08468 or 0.0309